OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, 9); Slope: 5 A. y – 1 = 5(x + 9) B. y – 9 = 5(x + 1) C. y – 9 = 5(x – 1) D. y + 9 = 5(x – 1)

OpenStudy (anonymous):

CAN SOMEBODY PLEASE HELP ME

OpenStudy (beccaboo333):

O_o I don't know >.< @thomaster can help. he's better with this stuff.

OpenStudy (beccaboo333):

@tHe_FiZiCx99 is good too.

OpenStudy (anonymous):

ok. Thanks. I need the answer. nut thanks anyways. @beccaboo333

OpenStudy (the_fizicx99):

You want the answer? The points are (1,9) and the slope is 5, right?

OpenStudy (anonymous):

yes. @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Well I can't give you the answer, put to find the point slope form, the want you to use y-y1=m(x-x1) x1 = 1 y1 9 M or slope = 5

OpenStudy (anonymous):

nevermind. I just need the answer right now. I dont even have time to try and work it out on here right now. I am sorry but I have to go if I cant just get the answer now

OpenStudy (the_fizicx99):

I'm not giving out the answers.

OpenStudy (the_fizicx99):

It's not even that hard. just plug in... gl.

OpenStudy (anonymous):

Ok. That is totes fine with me

OpenStudy (anonymous):

I got the answer. IT was c.